SOLVE BY USING QUADRATIC EQUATION
FIND THE CONSECUTIVE INTEGERS SUCH THAT THE SUM OF THEIR SQUARES IS 245
PLEASE HELP I HAD QUESTIONS LIKE THIS ON MY FINAL BUT I COULD NOT UNDER STAND HOW TO WORK THIS PROBLEM

Answer by mananth(16946)

let the number be x
the consequtive number will be x+1
sum of their squares = x^2 +(x+1)^2= 245
x^2 +(x+1)^2= 245
x^2+x^2+2x+1 =245
2x^2+2x-244=0
x^2+x-122=0
x^2 +12x-11x-122=0
x(x+12)-11(x+12)=0
(x+12(x-11)=0
x= -12 or 11
the numbers are -12 & -11 OR 11 & 12